# RetrievalAttention: Accelerating Long-Context LLM Inference via Vector Retrieval

### v2

@@ 3.1 Approximated Attention
Based on the above approximation, we build RetrievalAttention to only consider important key-value vectors (i.e., $\mathcal{I}_{t,\epsilon}$) which are retrieved by vector indexes.
@@ 3.2 Attention-aware Vector Search
For each pair of key and value vectors, we first decide whether to hold them in CPU or GPU memory (decision method is elaborated in §3.3). The KV vectors offloaded to CPU memory will be indexed by $\mathbf{k}_{i}\in\mathcal{R}^{d}$ and queried by $\mathbf{q}_{t}$ to find for the most relevant ones.
To accelerate the vector search during token generation, RetrievalAttention leverages the existing query vectors in the prefill phase to guide the index building for key vectors, efficiently mitigating the distribution gap. As shown in Figure 3(b), RetrievalAttention explicitly establishes connections from the query vector to its nearest key vectors (i.e., exact $k$-nearest neighbors, or KNN). The KNN results can be efficiently computed via GPU, forming a mapping from query vector distribution to key vector distribution. Using this structure, the decoding query vector can firstly searches its nearest query vectors and then obtain the most relevant key vectors through the distribution mapping.
Therefore, the KNN connections from query vectors to key vectors serves as a bridge to reconcile their distribution differences. However, this structure still has imperfections in both memory overhead and search efficiency because we need to store and access query vectors besides key vectors. To address this problem, we leverage the projection technique from the state-of-the-art cross-modal ANNS index RoarGraph [21] to eliminate all query vectors. Specifically, we project the KNN connections into key vectors by linking key vectors that are connected to the same query vectors, which efficiently streamlines the search. This process connects key vectors that are perceived close from query vectors’ perspective, allowing efficient index traversal for future query vectors.
Our evaluation shows that, by effectively modeling the proximity relationship between the query and key vectors, the vector database only requires scanning $1-3\%$ key vectors to reach a high recall, significantly reducing the index search latency by 74% compared with IVF indexes [14].
@@ 3.3 CPU-GPU Co-Execution
To exploit GPU parallelism and accelerate attention computation, RetrievalAttention decomposes the attention computation into two disjoint sets of KV cache vectors: the predictable ones on GPU and the dynamic ones on CPU, and then combine the partial attention outputs together.
We leverage the patterns observed in the prefill phase to predict KV vectors that are consistently activated during token generation. Similar to StreamingLLM [10], our current implementation uses fixed initial tokens and last sliding window of the context as the static pattern, and persist them in the GPU cache. RetrievalAttention can be adapted to utilize more complex static patterns [11, 22], achieving the best trade-off between low inference cost and high accuracy. To minimize data transfer over the slow PCIe interface, RetrievalAttention independently computes the attention results for the CPU and GPU components and then combines them, inspired by the FastAttention [23]. The detailed combination algorithm and overall execution flow of RetrievalAttention is illustrated in §B.
@@ 4 Evaluation
In this section, we compare the performance of RetrievalAttention in long-context LLM inference against full attention and other state-of-the-art methods. Through experiments, we mainly explore the following questions: (1) How does RetrievalAttention affect the model’s inference accuracy? Specifically, we assess the generation accuracy of RetrievalAttention and other methods across various downstream tasks. (§4.2) (2) Can RetrievalAttention efficiently reduce the token generation latency of long-context LLM inference? We compare the end-to-end decoding latency of RetrievalAttention with that of other baselines. (§4.3).
@@ 4.1 Experimental Setup
Testbed, Models, and Configurations. We conduct experiments on a server equipped with one NVIDIA RTX 4090 GPU (24GB memory) and an Intel i9-10900X CPU with 20 cores and 128GB DRAM. The experiment results using NVIDIA A100 GPU can be found in §A.4. We implement RetrievalAttention on three state-of-the-art long-context LLMs, including Llama-3-8B-Instruct-262k[24], Yi-6B-200K[25], and Yi-9B-200K[26]. To show practical speedup of RetrievalAttention and ensure the CPU memory consumption in long contexts do not exceed the capacity of DRAM, we follow previous work [5] to run the benchmark in real-world single-batch scenarios.
Baselines. We compare RetrievalAttention with the following training-free baselines. (1) Full attention without KV cache as well as the version with KV cache using vLLM [27]. (2) StreamingLLM[10]: it retains initial tokens along with fixed-length recent tokens in the GPU memory and discards remaining tokens. (3) SnapKV[11]: it caches the active tokens observed from the last window of the prompts. (4) InfLLM[6]: it separates KV cache of continuous token sequences into blocks and select representative vectors for each block. In the decoding phase, the current query scans all representative vectors and retrieve top-$k$ blocks with the highest similarity. We do not include Quest [5] for evaluation because its implementation  [^1] https://github.com/mit-han-lab/quest does not support the GQA model.
To better assess the effectiveness of our method, we introduce two additional baselines using traditional vector search methods from Faiss[16]. Specifically, Flat is an exact KNN method that performs a linear scan of all key-value vectors, whereas IVF indexes key vectors through clustering. By default, all indexing-based methods retrieve top-100 nearest key vectors to the current query vector.
@@
  • $\infty$-Bench[12]: this benchmark consists of 7 tasks, including three retrieval tasks (PassKey retrieval, Number retrieval, KV retrieval) and four realistic tasks (code debugging, dialogue and multiple-choices questions). The average context length of $\infty$-Bench is over 100K tokens.
  • RULER[13]: a comprehensive long-context benchmark consisting of 4 categories and 13 tasks, including retrieval, multi-hop tracing, aggregation, and QA tasks. It contains prompts of different lengths from 4K to 128K, allowing us to determine the actual context window size of models.
  • Needle-in-a-haystack [28]: it challenges the models to accurately retrieve information (the "needle") hidden within a lengthy document (the "haystack").
Table 2: Performance (%) of different methods and models on $\infty$-Bench. The size of the static pattern is consistently 640 (128 initial tokens + 512 tokens in the local window). All indexing based methods including Flat, IVF and RetrievalAttention retrieve top-100 key vectors by default. In the relatively complicated task KV Retrieval, we include the results of retrieving top-2000 key vectors.
Methods Tokens Retr.N Retr.P Retr.KV Code.D Math.F En.QA En.MC Avg.
Llama-3-8B FullAttention 128K 100.0 100.0 17.5 19.0 39.5 9.1 68.0 50.4
StreamingLLM 640 5.0 5.0 0.5 18.5 39.5 5.9 66.5 20.1 (-30.3)
@@ 4.2 Accuracy on Long Context Tasks
$\infty$-Bench. As shown in Table 2, RetrievalAttention achieves comparable accuracy to the full attention, benefiting from its efficient dynamic retrieval of important tokens. Static methods, such as StreamingLLM and SnapKV, lack this capability and therefore achieve sub-optimal accuracy. During token generation, the critical tokens change dynamically according to the current query, invalidating the previously captured static patterns. Although InfLLM supports dynamic retrieval of relevant blocks, it achieves nearly zero accuracy in complex tasks (i.e., KV retrieval) due to low accuracy of representative vectors. Since RetrievalAttention can accurately identify the most relevant key vectors, it achieves the best accuracy in KV retrieval. Moreover, by retrieving more tokens (i.e., top-2000 shown in the column of Retr.KV) in KV retrieval, RetrievalAttention achieves nearly the same accuracy as full attention, which demonstrating the effectiveness of our method in complex and dynamic tasks.
It is worth noting that Flat and IVF need to scan $100\%$ and $30\%$ of the past key vectors to achieve the same task accuracy as RetrievalAttention. In contrast, RetrievalAttention only requires scan $1-3\%$ vectors, resulting in much lower decoding latency.
RULER. Table 3 demonstrates that models utilizing RetrievalAttention achieves nearly the same task accuracy as full attention in different context lengths. In contrast, other training-free methods experience an noticeable reduction in accuracy, particularly for longer context sizes like 128K, as they fail to capture dynamically changed important tokens.
Needle-in-a-haystack. As shown in Figure 4, RetrievalAttention can effectively focus on information at various positions across different context windows, ranging from 4K to 128K. In contrast, methods like StreamingLLM encounter difficulties when critical information lies beyond the range of the static patterns, whose results are shown in §A.2.
4.3 Latency Evaluation
As the context length increases, the decoding latency of full attention significantly increases due to its quadratic time comlexity. Enabling the KV cache (vLLM) incurs out-of-memory (OOM) issues due to limited GPU memory. The latency of StreamingLLM, SnapKV, and InfLLM remains relatively stable because of constant tokens involved in the attention computation but they suffer significant accuracy degradation as demonstrated in §4.2. Due to efficient attention-aware vector search, RetrievalAttention achieves $4.9\times$ and $1.98\times$ latency reduction compared to Flat and IVF for the 128K context. The detailed latency breakdown can be seen in the Table 6 of §A.3.
@@ 4.3 Latency Evaluation
5 Related Works
To accelerate the long-context LLM inference, some works [29, 30, 10, 31, 32, 11] attempt to compress the size of the KV cache by leveraging the sparsity of attention. However, these methods often suffer from significant model accuracy drops due to the dynamic nature of attention sparsity.
@@ 5 Related Works
FlexGen [4] and Lamina [33] offload the KV cache to CPU memory, but they struggle with slow and costly full-attention computation. By identifying the dynamic nature of important KV vectors for different queries, recent work chooses to retain all of the KV cache and dynamically attend to different parts of KV vectors based on the current query. Quest [5] partitions the KV cache into blocks and selects a representative key vector for each block. For a given query, it scans all representative key vectors and attends top-$k$ blocks with the highest attention scores. InfLLM [6] adopts a similar strategy as Quest but offloads most KV cache blocks to the CPU memory to support longer contexts. Due to block-based organization and retrieval, the accuracy of representative vectors significantly impacts the effectiveness of those methods for obtaining important tokens. SparQ [7], InfiniGen [8], and LoKi [9] approximate the most relevant top-$k$ keys corresponding to a given query by reducing the head dimension. RetrievalAttention instead organizes the KV cache using ANNS indexes, allowing the retrieval of important tokens with high recalls and low cost. The concurrent work MagicPiG [34] and PQCache [35] employ locality-sensitive-hashing (LSH) and PQ centroids respectively to retrieve critical tokens. However, they fail to address the OOD issue in attention, necessitating retrieving a large portion of KV cache (e.g., 20%) to maintain the model accuracy.
@@ 6 Conclusion
We propose RetrievalAttention, a method that offloads most KV vectors to CPU memory and leverages vector search for dynamic sparse attention to minimize inference cost. RetrievalAttention identifies the different distributions of the query and key vectors and employs an attention-aware approach to efficiently find critical tokens for model generation. Experimental results demonstrate that RetrievalAttention effectively achieves $4.9\times$ and $1.98\times$ decoding speedup than exact KNN and traditional ANNS methods, on a single RTX4090 GPU for a context of 128K tokens. RetrievalAttention is the first system that supports running 8B-level LLMs with 128K tokens on a single 4090 (24GB) GPU with an acceptable latency cost and without compromising model accuracy.
@@ A.2 Additional Results on Needle-in-a-haystack
Table 6: Latency breakdown of retrieval attention-based algorithm on Llama-3-8B.
Methods Retrieval Attention Others Total
Flat 0.798 0.083 0.041 0.922
IVF 0.250 0.084 0.039 0.373
RetrievalAttention 0.064 0.081 0.043 0.188
A.3 Latency Breakdown
Table 6 presents the breakdown of end-to-end latency for different retrieval attention-based algorithms on a single RTX 4090 server under the 128K context length. Since RetrievalAttention effectively addresses the issue of out-of-distribution, it only requires 34.0% of the time for vector search. In contrast, Flat and IVF spends 86.6% and 67.0% of time, respectively. This is because RetrievalAttention scans less data, avoiding memory bandwidth contention when multiple heads are performing parallel retrieval in the CPU side. This advantage becomes more pronounced with longer context lengths.

### v3

@@ 3.1 Approximated Attention
Based on the above approximation, we build RetrievalAttention to only consider important key-value vectors (i.e., $\mathcal{I}_{t,\epsilon}$) that are persistent in GPU cache and dynamically retrieved by vector indexes.
@@ 3.2 Attention-aware Vector Search
For each pair of key and value vectors, we first decide whether to hold them in CPU or GPU memory (the decision method is elaborated in §3.3). The KV vectors offloaded to CPU memory will be indexed by $\mathbf{k}_{i}\in\mathcal{R}^{d}$ and queried by $\mathbf{q}_{t}$ to find the most relevant ones.
To accelerate the vector search during token generation, RetrievalAttention diverges from traditional indexes that only consider the closeness between key vectors for index building. Instead, it leverages the existing query vectors in the prefill phase to guide the index building for key vectors, efficiently mitigating the distribution gap. As shown in Figure 4(b), during the index construction, RetrievalAttention explicitly establishes connections from the query vector to its nearest key vectors (i.e., exact $k$-nearest neighbors, or KNN). The KNN results can be efficiently computed via GPU, forming a mapping from query vector distribution to key vector distribution. Using this structure, the decoding query vector can first search its nearest query vectors and then obtain the most relevant key vectors through the distribution mapping.
Therefore, the KNN connections from query vectors to key vectors serve as a bridge to reconcile their distribution differences. However, this structure still has imperfections in both memory overhead and search efficiency because we need to store and access query vectors besides key vectors. To address this problem, we leverage the projection technique from the state-of-the-art cross-modal ANNS index RoarGraph (Chen et al., 2024a) to eliminate all query vectors. Specifically, we project the KNN connections into key vectors by linking key vectors that are connected to the same query vectors, which efficiently streamlines the search. This process connects key vectors that are perceived as close from the query vectors’ perspective, allowing efficient index traversal for future query vectors.
Our evaluation shows that, by effectively modeling the proximity relationship between the query and key vectors, the vector database only requires scanning 1–3% key vectors to reach a high recall, significantly reducing the index search latency by 74% compared with IVF indexes (Sivic & Zisserman, 2003).
@@ 3.3 CPU-GPU Co-Execution
To exploit GPU parallelism and accelerate attention computation, RetrievalAttention decomposes the attention computation into two disjoint sets of KV cache vectors, the predictable ones on GPU and the dynamic ones on CPU, and then combines the partial attention outputs together.
We leverage the patterns observed in the prefill phase to predict KV vectors that are consistently activated during token generation. Similar to StreamingLLM (Xiao et al., 2024b), our current implementation uses fixed initial tokens and the last sliding window of the context as the static pattern and persists them in the GPU cache. RetrievalAttention can be adapted to utilize more complex static patterns (Li et al., 2024; Jiang et al., 2024), achieving the best trade-off between low inference cost and high accuracy. During the prefill phase, we physically separate the static tokens in the GPU memory with the remaining tokens, which are offloaded to the CPU memory indexed by the ANNS. To minimize data transfer over the slow PCIe interface, RetrievalAttention independently computes the attention results for the CPU and GPU components and then combines them, inspired by the FastAttention (Dao et al., 2022).
@@ 4 Evaluation
In this section, we compare the performance of RetrievalAttention in long-context LLM inference against full attention and other state-of-the-art methods. Through experiments, we mainly explore the following questions: (1) How does RetrievalAttention affect the model’s inference accuracy? Specifically, we assess the generation accuracy of RetrievalAttention and other methods across various downstream tasks (§4.2) (2) Can RetrievalAttention efficiently reduce the token generation latency of long-context LLM inference? We compare the end-to-end decoding latency of RetrievalAttention with that of other baselines (§4.3).
@@ 4.1 Experimental Setup
Testbed, Models, and Configurations. We conduct experiments on a server equipped with one NVIDIA RTX4090 GPU (24GB memory) and an Intel i9-10900X CPU with 10 physical cores (20 logical cores) and 128GB DRAM. The experiment results using NVIDIA A100 GPU can be found in §A.4. We implement RetrievalAttention on three state-of-the-art long-context LLMs, including Llama-3-8B-Instruct-262k (Gradient AI, 2024), Yi-6B-200K (01-ai, 2024a), and Yi-9B-200K (01-ai, 2024b). To show a practical speedup of RetrievalAttention and ensure the CPU memory consumption in long contexts does not exceed the DRAM capacity, we follow previous work (Tang et al., 2024) to run the benchmark in real-world single-batch scenarios.
Baselines. We compare RetrievalAttention with the following training-free baselines. (1) Full attention without KV cache as well as the version with KV cache using vLLM (Kwon et al., 2023). (2) StreamingLLM (Xiao et al., 2024b): it retains initial tokens along with fixed-length recent tokens in the GPU memory and discards remaining tokens. (3) SnapKV (Li et al., 2024): it only caches the critical tokens observed from the last window of the prompt. (4) InfLLM (Xiao et al., 2024a): it separates the KV cache of continuous token sequences into blocks and selects representative vectors for each block. In the decoding phase, the current query scans all representative vectors and retrieves top-$k$ blocks with the highest similarity. (5) Quest (Tang et al., 2024): it keeps track of the minimal and maximal key values in KV cache pages and estimates the criticality of a page using the query vector. (6) InfiniGen (Lee et al., 2024): it prefetches only the essential KV cache entries by speculating important tokens required for subsequent attention layers.
To better assess the effectiveness of our method, we introduce two additional baselines using traditional vector search methods from Faiss (Douze et al., 2024). Specifically, Flat is an exact KNN method that performs a linear scan of all key-value vectors, whereas IVF indexes key vectors through clustering. By default, all indexing-based methods retrieve the top-100 nearest key vectors.
@@
  • $\infty$-Bench (Zhang et al., 2024b): this benchmark consists of 7 tasks, including three retrieval tasks (passKey retrieval, number retrieval, KV retrieval) and four realistic tasks (code debugging, math find, dialogue and multiple-choices questions). The average context length of $\infty$-Bench is over 100K tokens.
  • RULER (Hsieh et al., 2024): a comprehensive long-context benchmark consisting of 4 categories and 13 tasks, including retrieval, multi-hop tracing, aggregation, and QA tasks. The prompt length ranges from 4K to 128K, allowing us to determine the actual context window size of models.
  • Needle-in-a-haystack (Greg Kamradt, 2023): it challenges the models to accurately retrieve information (the “needle”) hidden within a lengthy document (the “haystack”).
Table 2: Performance (%) of different methods and models on $\infty$-Bench. The size of the static pattern is consistently 640 (128 initial tokens + 512 tokens in the local window). All indexing-based methods, including Flat, IVF, and RetrievalAttention retrieve top-100 key vectors by default. In the relatively complicated task KV Retrieval, we include the results of retrieving top-2000 key vectors.
Methods Act. Tokens Retr.N Retr.P Retr.KV Code.D Math.F En.QA En.MC Avg.
Llama-3-8B
FullAttention 128K 100.0 100.0 17.5 19.0 39.5 9.1 68.0 50.4
StreamingLLM 2K 5.0 5.0 1.0 18.5 40.0 6.0 66.0 20.2 (-30.2)
@@ 4.2 Accuracy on Long Context Tasks
$\infty$-Bench. As shown in Table 2, RetrievalAttention achieves comparable accuracy to the full attention, benefiting from its efficient dynamic retrieval of important tokens. Static methods, such as StreamingLLM and SnapKV, lack this capability and, therefore, achieve sub-optimal accuracy. During token generation, the critical tokens change dynamically according to the current query, invalidating the previously captured static patterns. InfiniGen exhibits a noticeable drop in model accuracy compared to full attention due to inaccurate speculation of important tokens from previous layers. Although InfLLM and Quest supports dynamic retrieval of relevant blocks, it achieves nearly zero accuracy in complex tasks (i.e., KV retrieval) due to the low accuracy of representative vectors. Since RetrievalAttention can accurately identify the most relevant key vectors, it achieves the best accuracy in KV retrieval. Moreover, by retrieving more tokens (i.e., top-2000 shown in the column of Retr.KV) in KV retrieval, RetrievalAttention achieves nearly the same accuracy as full attention, which demonstrates the effectiveness of our method in complex and dynamic tasks.
It is worth noting that Flat and IVF need to scan 100% and 30% of the past key vectors to achieve the same task accuracy as RetrievalAttention. In contrast, RetrievalAttention only requires scan 1–3% vectors, resulting in much lower decoding latency.
RULER. Table 3 demonstrates that models utilizing RetrievalAttention achieve nearly the same task accuracy as full attention in different context lengths. In contrast, other training-free methods experience a noticeable reduction in accuracy, particularly for longer context sizes like 128K, as they fail to capture dynamically changed important tokens.
Needle-in-a-haystack. As shown in Figure 5, RetrievalAttention can effectively focus on information at various positions across different context windows, ranging from 4K to 128K. In contrast, other methods like StreamingLLM encounter difficulties when critical information lies beyond the range of the static patterns, whose results are shown in §A.2.
@@
Table 5: Decoding latency breakdown on Llama-3-8B.
Methods Retrieval Attention Others Total
Flat 0.798 0.083 0.041 0.922
IVF 0.250 0.084 0.039 0.373
RetrievalAttention 0.064 0.081 0.043 0.188
4.3 Decoding Latency
As the context length increases, the decoding latency of full attention significantly increases due to its quadratic time complexity. Enabling the KV cache (vLLM) incurs out-of-memory (OOM) issues due to limited GPU memory. The latency of StreamingLLM, SnapKV, and InfLLM remains relatively stable because of constant tokens involved in the attention computation, but they suffer significant model accuracy degradation. Due to efficient attention-aware vector search, RetrievalAttention achieves $4.9\times$ and $1.98\times$ latency reduction compared to Flat and IVF for the 128K context.
Table 5 presents the breakdown of end-to-end latency for different retrieval attention-based algorithms under the 128K context length. RetrievalAttention only requires 34.0% of the time for vector search, while Flat and IVF spend 86.6% and 67.0% of time, respectively. This is because RetrievalAttention scans less data for a high recall, avoiding memory bandwidth contention when multiple heads are performing parallel retrieval on the CPU side. Overall, compared with Flat and IVF, RetrievalAttention effectively reduces the index search latency by 91% and 74%, respectively. This advantage becomes more pronounced with longer context lengths.
Figure 6: Recall vs. scanning key vectors when using the query vector ($Q$ to $K$) and key vector ($K$ to $K$) as the query, individually. $Q$ and $K$ are dumped from three long-context LLM models.
4.4 Index Recall vs. Scanning Vectors
Now, we conduct a micro-analysis of the efficiency of attention-aware vector search by examining the relationship between recall and the number of scanned key vectors. The number of key vectors scanned to achieve a target recall serves as an indicator of search efficiency. Figure 6 demonstrates that for the $Q$ to $K$ search, RetrievalAttention requires scanning only a very limited number of key vectors (1–3%) to reach a recall rate higher than 0.95, whereas traditional indexes necessitate retrieving a significantly higher number of keys. We also included a well-known OOD-optimized solution RobustVamana (Jaiswal et al., 2022) for comparison. However, it performs poorly on attention vectors. The efficiency of RetrievalAttention arises because it effectively mitigates the OOD issue between query and key vectors. In contrast, for the in-distribution $K$ to $K$ search, all indexes exhibit good performance.
5 Related Works
To accelerate the long-context LLM inference, some works (Zhang et al., 2023; Liu et al., 2024b; Xiao et al., 2024b; Han et al., 2024; Ge et al., 2024; Li et al., 2024) attempt to compress the size of the KV cache by leveraging the sparsity of attention. However, these methods often suffer from significant model accuracy drops due to the dynamic nature of attention sparsity.
@@ 5 Related Works
FlexGen (Sheng et al., 2023) and Lamina (Chen et al., 2024b) offload the KV cache to CPU memory, but they struggle with slow and costly full-attention computation. By identifying the dynamic nature of important KV vectors for different queries, recent work chooses to retain all of the KV cache and dynamically attend to different parts of KV vectors based on the current query. Quest (Tang et al., 2024) partitions the KV cache into blocks and selects a representative key vector for each block. For a given query, it scans all representative key vectors and attends top-$k$ blocks with the highest attention scores. InfLLM (Xiao et al., 2024a) adopts a similar strategy as Quest but offloads most KV cache blocks to the CPU memory to support longer contexts. Due to block-based organization and retrieval, the accuracy of representative vectors significantly impacts the effectiveness of those methods for obtaining important tokens. SparQ (Ribar et al., 2024), InfiniGen (Lee et al., 2024), and LoKi (Singhania et al., 2024) approximate the most relevant top-$k$ keys corresponding to a given query by reducing the channel dimension. RetrievalAttention instead organizes the KV cache using ANNS indexes, allowing the retrieval of important tokens with high recalls and low cost. The concurrent work MagicPiG (Chen, 2024) and PQCache (Zhang et al., 2024a) employ LSH and PQ centroids to retrieve critical tokens, respectively. However, they fail to address the OOD issue in attention, necessitating retrieving a large portion of KV cache (e.g., 20%) for high model accuracy.
@@ 6 Conclusion
We propose RetrievalAttention, a method that offloads most KV vectors to CPU memory and leverages vector search for dynamic sparse attention to minimize inference cost. RetrievalAttention identifies the different distributions of the query and key vectors and employs an attention-aware approach to efficiently find critical tokens for model generation. Experimental results demonstrate that RetrievalAttention effectively achieves $4.9\times$ and $1.98\times$ decoding speedup than exact KNN and traditional ANNS methods, on a single RTX4090 GPU for a context of 128K tokens. RetrievalAttention is the first system that supports running 8B-level LLMs with 128K tokens on a single RTX4090 (24GB) GPU with an acceptable latency cost and without compromising model accuracy.
@@ A.2 Additional Results on Needle-in-a-haystack
A.3 Performance in the extremely long-context inference
Figure 8 shows the evaluation results of RetrievalAttention for extremely long contexts using the model Llama-3-8B-1048K. RetrievalAttention still passes all test cases when ranging the context length from 250K to 1 million, which demonstrates the robustness of our attention-aware indexes.
Figure 8: Performance of RetrievalAttention in 1 million Needle-in-a-haystack test.